OpenStudy (lgg23):

A particle (mass = 6.7 × 10^−27 kg, charge = 3.2 × 10^−19 C) moves along the positive x axis with a speed of 4.8 × 10^5 m/s. It enters a region of uniform electric field parallel to its motion and comes to rest after moving 2.0 m into the field. What is the magnitude of the electric field? Please explain and give formulas. I can do the math/work.

OpenStudy (badhi):

Kinetic energy stored in the particle = work done against the particle by the field $$\frac{1}{2} mv^2=F_ed=(eE)d\implies E=\frac{mv^2}{ed} $$

OpenStudy (mos1635):

2ed....

OpenStudy (mos1635):

E=mv^2 / 2ed

OpenStudy (badhi):

yeah, sorry for the typo

OpenStudy (lgg23):

Thank you very much!